Katie has a collection of nickels, dimes, and quarters with a total value of $4.65. There are 5 more dimes than nickels and 7 more quarters than nickels, how many of each coin does she have?

Answer by math_helper(2461)


Let
n = number of nickels
d = number of dimes
q = number of quarters
….
5n + 10d + 25q = 465 (total value, in cents)
d = n + 5 (5 more dimes than nickels)
q = n + 7 (7 more quarters than nickels)
The easiest thing to do at this point is to substitute for d & q in the top equation.
5n + 10(n+5) + 25(n+7) = 465
Can you solve for n?